If log10X is normally distributed with mean 4 and variance 4, find the probability of 1.202<X<83180000.(Given that log10 1202=3.08, log10 8310=3.92 )
1
Expert's answer
2020-04-20T05:34:53-0400

"P=P(1.202<X<83180000)"


let log normal distribution,

"P=P(\log_{10}1.202<\log_{10}X<\log_{10}83180000)"

"P=P(\log_{10}(\frac{1202}{1000}<\log_{10}X<\log_{10}(8318*10^4))\\
P=P(3.08-3<\log_{10}X<3.92+4)\\
P=P(0.08<\log_{10}X<7.92)\\"


convert to standard form using by

"z=\frac{x-\bar{x}}{\sigma}"

"P=P(\frac{0.08-4}{2}<Z<\frac{7.92-4}{2})\\
P=P(\frac{0.08-4}{2}<Z<\frac{7.92-4}{2})\\
P=P(-1.96<Z<1.96)\\
P=P(-1.96<Z<0)+P(0<Z<1.96)\\"

using calculator or table,

"P=0.475+0.475\\
\bold{P=0.95}"